OpenStudy (yrivers36):

Solve the given quadratic equation, if possible. 15n^2=7n+2 I set problem to 0 15n^2-7n+2=0 please help with explanation

OpenStudy (anonymous):

15n^2=7n+2 15n^2 - 7n - 2 = 0 (15 * -2 = -30 we split it as -10 * 3 as -10 + 3 = -7 ) 15n^2 - 10n + 3n - 2 = 0 5n(3n -2) + 1(3n - 2) = 0 (5n + 1)(3n -2) = 0 so either 5n + 1 = 0 or 3n - 2 = 0 if 5n + 1 = 0, then n = -1/5 if 3n - 2 = 0, then n = 2/3

OpenStudy (yrivers36):

Thank you Harkirat